OpenStudy (anonymous):

True or False: csc^-1[csc(-pi/4)]=-pi/4

hero (hero):

Hint: Use TI-Nspire Graphing Calc

OpenStudy (anonymous):

Do you know about TI - Nspire Graphing Calc @timmtamm96

hero (hero):

Most likely no

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

Now help him @Hero

OpenStudy (anonymous):

is there an online website for this calculator?

hero (hero):

Actually, there is.

OpenStudy (anonymous):

What may it be?

OpenStudy (anonymous):

Please and thank you :)

OpenStudy (anonymous):

\[\large cosec^{-1}(cosec (-\theta)) = - \theta\]

OpenStudy (anonymous):

I thought so. Thank you :)

OpenStudy (anonymous):

So what do you think is it true or false ??

OpenStudy (anonymous):

thrue

OpenStudy (anonymous):

You are also true then..

OpenStudy (anonymous):

true*

OpenStudy (anonymous):

Haha, thank you :)

OpenStudy (anonymous):

Welcome dear..

OpenStudy (anonymous):

true

OpenStudy (anonymous):

it is true csc^(-1)(csc(-pi/4)) = -pi/4